OpenStudy (anonymous):

Find the derivative of l(x)=4/x^2 by using the power rule. l ' (x) = _______ so would i start like this? 4x^2-1 ?

OpenStudy (helder_edwin):

actually \[\large \frac{4}{x^2}=4x^{-2} \] and apply the power rule for derivatives

OpenStudy (anonymous):

ohh okay, darn.. :P so would it be like this? (-1) * 4x^-2 = -4x^-2 = -4/x^2 ? :/

OpenStudy (agent0smith):

Multiply by the power (which is -2). Then, reduce the power by one.

OpenStudy (anonymous):

ohh so (-1) * 4x^2-1 = (-1) * 4x^1 = -4x ?

OpenStudy (agent0smith):

No... Read the above again :P

OpenStudy (agent0smith):

And the power is -2 NOT 2.

OpenStudy (anonymous):

ohhh okay oops.. :P so (-1) * 4x^-2-1=(-1) * 4x^-3 = -4x^-3 = -4/x^3 ??

OpenStudy (agent0smith):

Start from scratch. \[\large I(x) = 4x^{-2}\] Multiply by the power (which is -2). Then, reduce the power by one.

OpenStudy (anonymous):

=4x^-2-1= 4x^-3 ?

OpenStudy (agent0smith):

Read the power rule again... you're skipping an entire step.

OpenStudy (helder_edwin):

u have \[\large f(x)=Kx^n \] where K=4 and n=-2 the power rule says \[\large f'(x)=Knx^{n-1} \]

OpenStudy (anonymous):

oh so (4) x (-2^-2-1) ?

OpenStudy (agent0smith):

^ can't tell what you are doing... \[\large I(x) = 4x^{-2}\] Multiply by the power (which is -2). Then, reduce the power by one. \[\large I'(x) = -2* 4x^{-2-1}\]

OpenStudy (anonymous):

ohh okay and sorrry!! i'll try to make it neater!! and more eligible lol :P so -2 * 4x^ -3 =-8x^-3 = -8/x^3 ??

OpenStudy (agent0smith):

Use the equation editor, it's not too hard. You can right click what i write and copy and paste

OpenStudy (helder_edwin):

yes. u got it.

OpenStudy (anonymous):

\[I′(x)=−2∗4x−2−1 \] \[= -8/x^3\]

OpenStudy (anonymous):

haha whoah.. never used equation tool before!! it's like latex!! haha

OpenStudy (anonymous):

oh yayy!! :)